OpenStudy (anonymous):

Whoever can figure this out and explain is so awesome! You saw a shirt for $97. You didn’t have the cash, so you borrowed $50 from your mom and $50 from your dad = $100. You bought the shirt, and had $3 change. You gave your dad $1 and your mom $1 and kept the other $1 for yourself. Now you owe your mom $49 and your dad $49. 49+49 = 98 + your $1 = 99. Where is the missing $1?

OpenStudy (anonymous):

Tax? lol

OpenStudy (anonymous):

You don't owe BOTH parents $49. You owe ONE OF THEM $49, whilst the other only has $48, COZ YOU TOOK THE REMAINING $1. Initially, ONE PARENT owes you $49, and the other owes your $48 ($1 to one parent, $2 to the other OR $1.5 to both but this is more complicated lol). But since YOU TOOK the remaining $1... well? Can you see it now???? Because it is an ODD number, you get an ODD number for a change. HENCE. ONE parent gets MORE than the other in return, but since you TOOK the dollar, the rule still applies. You are still missing 1 dollar. $49 + $48 = $97 ($1 given to mum, dad, and you took one) IF you get back from yo parents the $1 you gave them: (49+1) + (48+1) = $99 The remaining dollar is the one that you took! If you DIDN'T take the dollar, you would have given BOTH parents $1.50 each (1.5x2 = $3) It would still add up to $100.

Parth (parthkohli):

^ COPY/PASTE

OpenStudy (anonymous):

yessss

Parth (parthkohli):

I know becky can't type so fast ;)

OpenStudy (anonymous):

hahahahhaha, good one xD

OpenStudy (anonymous):

Well actually the typing I could do...the problem however...

OpenStudy (anonymous):

"You gave your dad $1 and your mom $1" so you would still owe both $49. There were $3 extra.

Parth (parthkohli):

Yes exactly

OpenStudy (anonymous):

you said 49 + 49 = 98 this is what you still owe to mom and dad. why, because you took $100, paid $97 for the shirt and kept $1 for yourself. so this $1 in question is a part of the money that you owe to your parents. i.e, part of the $98 you owe to your parents. So you do not add it to the $98, because it is already a part of the $98, which you owe. suppose you give the $1 to your parents , now, you still owe then $98 - $1, i.e, $97(which is the money you actually spent on the shirt hah this one?

OpenStudy (anonymous):

the remaining $1 is in the payment for the shirt

OpenStudy (anonymous):

haha Idk this is to complicated for my brain.

OpenStudy (anonymous):

Confession! I kinda, sorta, googled it :3

Parth (parthkohli):

Lol I have read this problem I was letting the ones who were typing do it :D

OpenStudy (anonymous):

haha I do that all the time :p

OpenStudy (anonymous):

The confusion comes in the wording. Let me simply explain the situation in a simpler, less confusing manner. You borrowed 100 dollars, spent 97 and got 3 dollars change. You used 2 dollars of that to start paying your parents back, but you still need 98 more dollars to pay them back. It turns out you only have 1 left dollar left, so you need 97 more dollars to finish paying them. In short: Money borrowed = 100 2 dollars returned, 1 dollar you have, 97 more that you need to find somehow. 2 + 1 + 97 = 100. It works out.

OpenStudy (anonymous):

In particular, this is the sentence that is stupid. Now you owe your mom $49 and your dad $49. 49+49 = 98 + your $1 = 99. Where is the missing $1? It should be 49+49 =98 +the $2 you already returned = 100

OpenStudy (anonymous):

Taking the numbers out of it: Money borrowed = (Money already returned) + (Money still needing to be returned) Instead, they wrote: Money borrowed = (Money already returned) + (Money that I HAVE) which is stupid, because I could HAVE any amount of money.

OpenStudy (anonymous):

Thanks so much me and my grandpa were going crazy :p

OpenStudy (anonymous):

My pleasure =D